Solve y2 + 8y = -15 by factoring.

Answer by algebrahouse.com(1659)

"Solve y2 + 8y = -15 by factoring."

y² + 8y = -15
y² + 8y + 15 = 0 {added 15 to both sides}
(y + 5)(y + 3) = 0 {factored into two binomials}
y + 5 = 0 or y + 3 = 0 {set each factor equal to 0}
y = -5 or y = -3 {solved each equation for x}
www.algebrahouse.com